OpenStudy (anonymous):

The product of (a + b)(a − b) is a^2 − b^2

OpenStudy (anonymous):

=a^2-ab+ab-b^2=a^2-b^2

OpenStudy (anonymous):

Answers: sometimes always or never

OpenStudy (anonymous):

always